### v1

@@ 5.1. Continuous group cohomology and $A_{3}$-formality
Let $G$ be a profinite group, and let $G^{n}$ denote the $n$-fold direct product of $G$ with itself. Let $p$ be a prime number. Let $\mathcal{C}^{n}(G,\mathbb{F}_{p})$ denote the $\mathbb{F}_{p}$-vector space of continuous functions $G^{n}\to\mathbb{F}_{p}$ with respect to the discrete topology on $\mathbb{F}_{p}$ and the profinite topology on $G$. Following [42, §2.2] the differential $\delta\colon\mathcal{C}^{n}(G,\mathbb{F}_{p})\to\mathcal{C}^{n+1}(G,\mathbb{F}_{p})$, which is defined by
@@
Free pro-$p$ groups satisfy the following much stronger property: A graded $\mathbb{F}_{p}$-algebra $A$ is called intrinsically $A_{\infty}$-formal if every differential graded algebra $\mathcal{A}$ with $H^{\bullet}(\mathcal{A})\cong A$ is $A_{\infty}$-formal.
@@
Since the inclusion map of the Koszul complex into the bar complex is the identity, $H^{\bullet}(G,\mathbb{F}_{p})$ is Koszul. Hence we may compute the Hochschild cohomology of $H^{\bullet}(G,\mathbb{F}_{p})$ using the Koszul complex. Since $H^{i}(G,\mathbb{F}_{p})$ is trivial for $i\geq 2$, the groups $\mathrm{HH}^{n,s}(H^{\bullet}(G,\mathbb{F}_{p}))$ for $n+s\geq 2$ are trivial. In particular, we have $\mathrm{HH}^{n,2-n}(H^{\bullet}(G,\mathbb{F}_{p}))$ for all $n\geq 3$. By Kadeishvili’s theorem, proven also in [40, Theorem 4.7, page 85], this implies that $\mathcal{C}^{\bullet}(G,\mathbb{F}_{p})$ is intrinsically $A_{\infty}$-formal. ∎
@@
Recall that the only finite Demushkin group is $G=\mathbb{Z}/2\mathbb{Z}$. In this case, $H^{\bullet}(G,\mathbb{F}_{2})$ is isomorphic to the polynomial algebra $\mathbb{F}_{2}[x]$ in one generator. The latter is a Koszul algebra with no relations and hence $K_{n}^{n}(H^{\bullet}(G,\mathbb{F}_{2}))$ is trivial for all $n\geq 2$. As in the proof of Proposition 5.3, this implies that $H^{\bullet}(G,\mathbb{F}_{2})$ is intrinsically formal. See also [34, Lemma 6.2].
@@ 5.2. Dwyer’s criterion
We now recall from [5, Theorem 2.4] that the vanishing of triple Massey products in group cohomology can be characterised as follows. Let $U_{n}(\mathbb{F}_{p})$ denote the group of all upper triangular unipotent $(n\times n)$-matrices with coefficients in $\mathbb{F}_{p}$. Let $Z_{n}(\mathbb{F}_{p})$ denote the center of $U_{n}(\mathbb{F}_{p})$, i.e., the subgroup of all matrices in $U_{n}(\mathbb{F}_{p})$ with all off-diagonal entries being 0 except at position $(1,n)$. Write $\overline{U}_{n}(\mathbb{F}_{p})=U_{n}(\mathbb{F}_{p})/Z_{n}(\mathbb{F}_{p})$.
@@
The following result is a special case of [5, Theorem 2.4].
@@ 5.3. Demushkin groups are Koszul
We now show that the cohomology algebra of a Demushkin group is $A_{3}$-formal. We will deduce this fact from a more general result on quadratic algebras which is well known and proven for example in [36, Proposition 2.3 in Chapter 2 on page 24, and Proposition 4.2 of Chapter 6 on page 124]. We provide a proof for completeness and convenience of the reader.
@@
is exact. Thus, $A\otimes K_{\bullet}^{\bullet}(A)$ provides a free resolution of $\mathbb{F}$. This implies that $K(A)$ is a minimal free resolution of $A$ as an $A$-bimodule by [43, Proposition 3.1], where we note that, by [1, Section 2.8], the complex $K^{\prime}(A)$ in [43] is isomorphic to the complex we denote by $K(A)$. The two-out-of-three property of quasi-isomorphisms then implies that $A$ is Koszul as defined in Definition 4.3
@@
Corollary 5.12 also follows from [32, Theorem 5.2] where a stronger result is proven. The proof of [32, Theorem 5.2] uses the fact that the dual algebra $(H^{\bullet})^{!}=T((H^{1})^{*})/(R^{\perp})$ of $H^{\bullet}=H^{\bullet}(G,\mathbb{F}_{p})$ is a quadratic algebra which satisfies the assumption of [32, Lemma 2.15], i.e., $\chi_{1}^{*}\otimes\chi_{1}^{*}\notin R^{\perp}$ while $\chi^{*}_{1}\otimes\chi_{2}^{*}\in R^{\perp}$, where $\chi_{i}^{*}$ denotes basis vectors of $(H^{1})^{*}$ which are dual to the $\chi_{i}$. This implies that $(H^{\bullet})^{!}$ has a Poincaré–Birkhoff–Witt basis which is known to imply that an algebra is Koszul. Then it remains to use the general fact that a locally finite quadratic algebra is Koszul if and only if its quadratic dual is Koszul.
@@ 6.2. The canonical class of a Demushkin group
Let $\mathcal{C}^{\bullet}=\mathcal{C}^{\bullet}(G,\mathbb{F}_{p})$ denote the complex of continuous inhomogeneous cochains, let $\mathcal{Z}^{\bullet}=\mathcal{Z}^{\bullet}(G,\mathbb{F}_{p})$ denote the cocycles, and let $H^{\bullet}=H^{\bullet}(G,\mathbb{F}_{p})$ denote the corresponding cohomology algebra. Let $\{\chi_{1},\ldots,\chi_{d}\}$ be an $\mathbb{F}_{p}$-basis of $H^{1}$ such that $\chi_{i}(x_{j})=-\delta_{ij}$, where $\delta_{ij}$ denotes the Kronecker symbol. The $\mathbb{F}_{p}$-vector space $H^{2}$ is then generated by the single element $\chi_{1}\cup\chi_{2}$. We have $H^{\bullet}=T(H^{1})/(R)$ where, by [33, Proposition 3.9.13] (see also [32, page 27]), $R\subset H^{1}\otimes H^{1}$ is the $\mathbb{F}_{p}$-vector subspace $R=\mathrm{span}(\mathbf{B})$ where $\mathbf{B}$ is the set
@@ 6.3. Demushkin groups with invariant $q=3$ are not $A_{3}$-formal
We now provide an alternative proof of Theorem 6.12. Let $G$ be the pro-$3$ group as in Theorem 6.12. It is well-known that triple Massey products vanish for Demushkin groups (see for example [31, Theorem 4.3] and [35, Theorem 3.5]). We therefore point out that the argument in the proof of Theorem 6.12 does not imply that the triple Massey product $\langle\chi_{1},\chi_{1},\chi_{1}\rangle$ does not vanish. In fact, the argument shows that for the particular defining system $(\chi_{1},\chi_{1},\chi_{1},-f_{2}(\chi_{1}\otimes\chi_{1}),-f_{2}(\chi_{1}\otimes\chi_{1}))$, the cocycle
@@
By [19, page 254], the group $G$ of Theorem 6.12 for $d=4$ is realisable as the maximal pro-$3$ Galois group $G_{F}(3)$ of the field $F=\mathbb{Q}_{3}(\zeta_{3})$ where $\zeta_{3}$ is a root of unity of order $3$. According to [6] and [19] (see also [39, Remark 3.3]), this is the only Demushkin group of rank $4$ which is known to be realisable as the maximal pro-$p$ Galois group $G_{F}(p)$ of a field.

### v2

@@ 5.1. Continuous group cohomology and $A_{3}$-formality
Let $G$ be a profinite group, and let $G^{n}$ denote the $n$-fold direct product of $G$ with itself. Let $p$ be a prime number. Let $\mathcal{C}^{n}(G,\mathbb{F}_{p})$ denote the $\mathbb{F}_{p}$-vector space of continuous functions $G^{n}\to\mathbb{F}_{p}$ with respect to the discrete topology on $\mathbb{F}_{p}$ and the profinite topology on $G$. Following [48, §2.2] the differential $\delta\colon\mathcal{C}^{n}(G,\mathbb{F}_{p})\to\mathcal{C}^{n+1}(G,\mathbb{F}_{p})$, which is defined by
@@
Free pro-$p$ groups satisfy the following much stronger property (see also Remark 3.12): A graded $\mathbb{F}_{p}$-algebra $A$ is called intrinsically $A_{\infty}$-formal if every differential graded algebra $\mathcal{A}$ with $H^{\bullet}(\mathcal{A})\cong A$ is $A_{\infty}$-formal.
@@
Since the inclusion map of the Koszul complex into the bar complex is the identity, $H^{\bullet}(G,\mathbb{F}_{p})$ is Koszul. Hence we may compute the Hochschild cohomology of $H^{\bullet}(G,\mathbb{F}_{p})$ using the Koszul complex. Since $H^{i}(G,\mathbb{F}_{p})$ is trivial for $i\geq 2$, the groups $\mathrm{HH}^{n,s}(H^{\bullet}(G,\mathbb{F}_{p}))$ for $n+s\geq 2$ are trivial. In particular, we have $\mathrm{HH}^{n,2-n}(H^{\bullet}(G,\mathbb{F}_{p}))$ for all $n\geq 3$. By Kadeishvili’s theorem, proven also in [46, Theorem 4.7, page 85], this implies that $\mathcal{C}^{\bullet}(G,\mathbb{F}_{p})$ is intrinsically $A_{\infty}$-formal. ∎
@@
Recall that the only finite Demushkin group is $G=\mathbb{Z}/2\mathbb{Z}$. In this case, $H^{\bullet}(G,\mathbb{F}_{2})$ is isomorphic to the polynomial algebra $\mathbb{F}_{2}[x]$ in one generator. The latter is a Koszul algebra with no relations and hence $K_{n}^{n}(H^{\bullet}(G,\mathbb{F}_{2}))$ is trivial for all $n\geq 2$. As in the proof of Proposition 5.3, this implies that $H^{\bullet}(G,\mathbb{F}_{2})$ is intrinsically formal. See also [39, Lemma 6.2].
@@ 5.2. Dwyer’s criterion
We now recall from [6, Theorem 2.4] that the vanishing of triple Massey products in group cohomology can be characterised as follows. Let $U_{n}(\mathbb{F}_{p})$ denote the group of all upper triangular unipotent $(n\times n)$-matrices with coefficients in $\mathbb{F}_{p}$. Let $Z_{n}(\mathbb{F}_{p})$ denote the center of $U_{n}(\mathbb{F}_{p})$, i.e., the subgroup of all matrices in $U_{n}(\mathbb{F}_{p})$ with all off-diagonal entries being 0 except at position $(1,n)$. Write $\overline{U}_{n}(\mathbb{F}_{p})=U_{n}(\mathbb{F}_{p})/Z_{n}(\mathbb{F}_{p})$.
@@
The following result is a special case of [6, Theorem 2.4].
@@ 5.3. Demushkin groups are Koszul
We now show that the cohomology algebra of a Demushkin group is $A_{3}$-formal. We will deduce this fact from a more general result on quadratic algebras which is well known and proven for example in [41, Proposition 2.3 in Chapter 2 on page 24, and Proposition 4.2 of Chapter 6 on page 124]. We provide a proof for completeness and convenience of the reader.
@@
is exact. Thus, $A\otimes K_{\bullet}^{\bullet}(A)$ provides a free resolution of $\mathbb{F}$. This implies that $K(A)$ is a minimal free resolution of $A$ as an $A$-bimodule by [49, Proposition 3.1], where we note that, by [1, Section 2.8], the complex $K^{\prime}(A)$ in [49] is isomorphic to the complex we denote by $K(A)$. The two-out-of-three property of quasi-isomorphisms then implies that $A$ is Koszul as defined in Definition 4.3
@@
Corollary 5.12 also follows from [31, Theorem 5.2] where a stronger result is proven. The proof of [31, Theorem 5.2] uses the fact that the dual algebra $(H^{\bullet})^{!}=T((H^{1})^{*})/(R^{\perp})$ of $H^{\bullet}=H^{\bullet}(G,\mathbb{F}_{p})$ is a quadratic algebra which satisfies the assumption of [31, Lemma 2.15], i.e., $\chi_{1}^{*}\otimes\chi_{1}^{*}\notin R^{\perp}$ while $\chi^{*}_{1}\otimes\chi_{2}^{*}\in R^{\perp}$, where $\chi_{i}^{*}$ denotes basis vectors of $(H^{1})^{*}$ which are dual to the $\chi_{i}$. This implies that $(H^{\bullet})^{!}$ has a Poincaré–Birkhoff–Witt basis which is known to imply that an algebra is Koszul. Then it remains to use the general fact that a locally finite quadratic algebra is Koszul if and only if its quadratic dual is Koszul.
@@ 6.2. The canonical class of a Demushkin group
Let $\mathcal{C}^{\bullet}=\mathcal{C}^{\bullet}(G,\mathbb{F}_{p})$ denote the complex of continuous inhomogeneous cochains, let $\mathcal{Z}^{\bullet}=\mathcal{Z}^{\bullet}(G,\mathbb{F}_{p})$ denote the cocycles, and let $H^{\bullet}=H^{\bullet}(G,\mathbb{F}_{p})$ denote the corresponding cohomology algebra. Let $\{\chi_{1},\ldots,\chi_{d}\}$ be an $\mathbb{F}_{p}$-basis of $H^{1}$ such that $\chi_{i}(x_{j})=-\delta_{ij}$, where $\delta_{ij}$ denotes the Kronecker symbol. The $\mathbb{F}_{p}$-vector space $H^{2}$ is then generated by the single element $\chi_{1}\cup\chi_{2}$. We have $H^{\bullet}=T(H^{1})/(R)$ where, by [38, Proposition 3.9.13] (see also [31, page 27]), $R\subset H^{1}\otimes H^{1}$ is the $\mathbb{F}_{p}$-vector subspace $R=\mathrm{span}(\mathbf{B})$ where $\mathbf{B}$ is the set
@@ 6.3. Demushkin groups with invariant $q=3$ are not $A_{3}$-formal
We now provide an alternative proof of Theorem 6.12. Let $G$ be the pro-$3$ group as in Theorem 6.12. It is well-known that triple Massey products vanish for Demushkin groups (see for example [36, Theorem 4.3] and [40, Theorem 3.5]). We therefore point out that the argument in the proof of Theorem 6.12 does not imply that the triple Massey product $\langle\chi_{1},\chi_{1},\chi_{1}\rangle$ does not vanish. In fact, the argument shows that for the particular defining system $(\chi_{1},\chi_{1},\chi_{1},-f_{2}(\chi_{1}\otimes\chi_{1}),-f_{2}(\chi_{1}\otimes\chi_{1}))$, the cocycle
@@
By [20, page 254], the group $G$ of Theorem 6.12 for $d=4$ is realisable as the maximal pro-$3$ Galois group $G_{F}(3)$ of the field $F=\mathbb{Q}_{3}(\zeta_{3})$ where $\zeta_{3}$ is a root of unity of order $3$. According to [7] and [20] (see also [45, Remark 3.3]), this is the only Demushkin group of rank $4$ which is known to be realisable as the maximal pro-$p$ Galois group $G_{F}(p)$ of a field.